OpenStudy (anonymous):

The population of a town is 5000.if the males increased by 6% and the females by 14% the population will be 5500.find the number of females in the town

OpenStudy (anonymous):

number of females is 2500

OpenStudy (asevilla5):

http://openstudy.com/study#/updates/4de7514b4c0e8b0b0d84b8d8 here is a question very similar might help

OpenStudy (anonymous):

Let number of males be m and number of females be f. Therefore, m + f = 5000 -(1) Increased by 6% and 14% means +0.06 and +0.14 respectively, Therefore, 1.06m + 1.14f = 5500 -(2) Continue and solve from here